Speed of light in a transparent medium is 2x108m/s. What is the index of refraction for this medium? c=3x108m/s

a.

0.66

b.

1

c.

1.5

d.

2
Answer #1

here,

speed of light in transparent medium , v = 2 * 10^8 m/s

let the index of refraction be n1

using snell's law

n1/n2 = c/v

n1 / 1 = 3/2

n1 = 1.5

the index of refraction of medium is 1.5 m/s